OpenStudy (anonymous):

Devin wants to have equal amounts of room on all four sides of his ping pong table. A 5ft by 9ft table is placed in the center of a rectangular room that has an area of 221 square feet. What is the width of the clear space Devin will have on all sides of his table? Show your work.

OpenStudy (whpalmer4):

It sounds like we get to choose the dimensions of the room, so long as the area is 221 ft^2 and it gives Devin equal borders around his ping-pong table. Do you agree with that interpretation?

OpenStudy (anonymous):

YES

OpenStudy (whpalmer4):

Okay, and we'll assume that when they say "equal amounts of room" that really means the distance from the table to the wall will be the same on each side, not that the area will be the same.  So the length of the room will be 9 + x + x, and the width of the room will be 5 + x + x, right?

OpenStudy (whpalmer4):

Remember the formula for the area is A = l * w, and the A = 221. If l = 9+x+x and w = 5+x+x, can you write the equation for A in terms of x?